Please Help with this problem: You have just won 100,000.00 from a lottery. If you invest all this amount in a tax-free money market fund earning 7% compounded weekly, how long do you have to wait to become a millionaire?: Please Help with this problem: You have just won 100,000.00 from a lottery. If you invest all this amount in a tax-free money market fund earning 7% compounded weekly, how long do you have to wait to become a millionaire?
Answer by ankor@dixie-net.com(4189)

100,000.00 from a lottery. If you invest all this amount in a tax-free money market fund earning 7% compounded weekly, how long do you have to wait to become a millionaire?
:
P(1+(r/n))^t = A
Where
P = 100000
A = 1000000
r = .07
n = 52
t = time in weeks
:
100000(1+(.07/52)) = 1000000
:
1.0013461538)^t = 10; divide both sides by 100000
:
t * ln(1.001346154) = ln(10); find the natural log of both sides
:
.0013452487t = 2.302585
t = 2.302585/.0013452487
t = 1711.64 ~ 1712 weeks
or
1712/52 ~ 33 years
:
:
Check solution on calc: enter 100000(1+(.07/52))^1711.64 = 99999.62 ~ 1000000